{3v-2w=1
{2v+2w=4

Answer by mananth(16949)

3v-2w=1
2v+2w=4
...
add the two equations
3v-2w+2v+2w=1+4
5v=5
/5
v=1
...
plug value of v in second equation
2v+2w=4
2*1+2w=4
2w=4-2
2w=2
/2
w=1
...
v=1 ,w=1
m.ananth@hotmail.ca